# TermSight: Making Service Contracts Approachable

### v1

@@ 6.4. Analysis
In this section, we introduce our analysis framework for the user study. We used a causal framework (§6.4.1) to understand the effects of the treatment and a Bayesian analysis (§6.4.2) to estimate the treatment effect. Then, we discuss the thematic analysis (§6.4.3) used to analyze the qualitative responses from the semi-structured interviews.
6.4.1. Structural Causal Framework
We used a structural causal framework popularized by Pearl (2009) to better understand the effects of the treatment. A structural causal model framework comprises exogenous variables $U$, endogenous variables $V$ and a set of functions $F$ that represent the physical process that results in the values of the endogenous variables. For example, the question $v_{i}=f(v,u)$ implies that the values of $v$ and $u$ completely determine the value of $v_{i}$ through the function $f$. The structural causal model can be represented as a Directed Acyclic Graph (DAG) where the nodes represent the variables and the edges represent the causal relationships among the variables. Thus, in the preceding example, $v$ and $u$ are the parents of $v_{i}$ in the DAG. The structural causal model framework is non-parametric, but frequently used with generalized linear models to estimate the functions $f$. The structural causal framework has connections with structural equation models (Wright, 1934) used in the Economic Sciences, though the latter were developed with linear relationships in mind. Figure 9 shows the DAG for the study.
[…]
Figure 9. The causal graph (a DAG) has three types of nodes: Treatment ($T$) colored blue, pre-treatment covariates $D$ (demographics), $K$ (knowledge of Terms of Service) colored gray, and measured outcomes $Q$ (comprehension), $R$ (recall of correct facts), and $E$ (experience of the interface) colored orange. Since we use randomized assignment, including the interface type (treatment vs. control), type of ToS (social media vs. e-commerce website), and order of presentation (first vs. second), the treatment $T$ is independent of the pre-treatment covariates $D$ and $K$. Thus, to estimate the causal effect of the treatment $T$ on the outcomes $Q$, $E$, and $R$, we can use a simple regression model, without conditioning on the pre-treatment covariates $D$ and $K$ since there is no backdoor path from the treatment $T$ to the outcome variables $Q$, $E$, and $R$.
@@ 6.4.2. Bayesian Analysis
We used Bayesian inference to estimate the treatment effect of TermSight on the outcomes of interest. While the use of Bayesian estimates is growing in HCI (Kay et al., 2016; Koshy et al., 2023), we briefly justify its use over traditional non-Bayesian methods. Standard statistical methods including t-tests of significance can be powerful, but they require us to test for the conditions under which a test is valid (e.g., Normality for the t-test). Prior work in Biological and Psychological Sciences has shown that Null Hypothesis Statistical Testing (NHST) including $p$-values are problematic, and must be used with caution. They depend on the researchers’ intentions, when they stop collecting samples, and pre-selection of hypothesis (Ioannidis, 2005, 2012; Button et al., 2013).
We believe that Bayesian techniques offer several advantages over non-Bayesian methods. First, as Kay et al. (2016) point out, a Bayesian framework leads to an accumulation of knowledge within HCI, where the posterior of the parameters in a prior work can serve as the prior in the current experiment. Second, a Bayesian model is transparent—the researcher will foreground all the assumptions in the analysis via their model, allowing for a principled critique of the model. Third, use of a Bayesian framework shifts the discussion from “did it work” to “effect size of the intervention” (Kay et al., 2016). While NHST techniques can compute the confidence interval, these intervals are susceptible to misinterpretation (Belia et al., 2005; Hoekstra et al., 2014) and importantly, underemphasized in favor of $p$-values. Finally, as  McElreath (2015) points out, a Bayesian model with use of maximum entropy priors for the parameters (e.g., the normal distribution) is the most conservative given the evidence to estimate of the effect of the treatment.
@@ 7.1. RQ1: How did participants perceive TermSight and its features?
When asked about which version of the interface they prefer, 19 participants preferred TermSight. One participant preferred the baseline interface, citing difficulty navigating TermSight’s two scrollable columns on a small screen with a trackpad that lacks scroll functionality. Below we present our quantitative findings on the experience outcomes (§7.1.1), followed by qualitative findings organized around recurring themes about features of TermSight.
Figure 10. 5-Point Likert Ratings of participants’ reading experiences: Participants found it to be easier to find and understand relevant information. Participants were also more willing to spend more time on the ToS with TermSight and read other ToS with TermSight. The 7 experience questions are specified in Appendix F.3.
7.1.1. Quantitative Findings: TermSight improved user experience measures
@@
Figure 10 shows participants’ ratings of the 7 reading experience measures for both interfaces. Our Bayesian model analysis (model details in Appendix D.1) reveals a significant effect of the TermSight interface on the reading experience measures. The forest plots shown in Figure 19 in the Appendix demonstrate that the 94% Highest Posterior Density Interval (HPDI [^1] In Bayesian analysis, the 94% HPDI refers to that smallest interval (this interval is unique) of the posterior distribution that has 94% of the probability mass. It is common in Bayesian analysis to use intervals distinct (97% and 89% HPDI intervals are also used) from the typical 95% value to avoid the confusion of the frequentist confidence intervals.) for each treatment–control pair for every user experience measure does not overlap with each other. This indicates a significant positive effect of the treatment on all 7 user experience measures.
In Figure 11, we show the contrast between the treatment and control conditions averaged across 7 user experience measures as well as the posterior of the effect size. The posterior distribution (Figure 11(b)) of the effect size (Cohen’s $d$) of the treatment on the user experience is centered around 2.7, with an HPDI of [2.3, 3.2], indicating a large effect size [^2] This is the effect size measured not on the outcome Likert scale (1–5) but on latent scale of the model coefficients.. Because the HPDI doesn’t overlap with the ROPE, the effect of TermSight on user experience is significant. Below, we further present recurring themes related to participants’ perceptions and feedback on features of TermSight.
(a) Posterior distributions of the coefficients for the treatment and control conditions.
(b) Posterior distribution of the effect size of the treatment on the user experience.
@@
Figure 11. Figure 11(a) shows the posterior distributions of the coefficients for the treatment and control conditions. Note that the posterior distribution of the treatment condition is shifted to the right, indicating a positive effect of the treatment on the user experience. Figure 11(b) shows the posterior distribution of the effect size of the treatment on the user experience centered around 2.7, suggesting a large effect size. Additionally, the HPDI doesn’t overlap with the ROPE, suggesting a significant effect of the treatment on user experience.
7.1.2. Power Meter provided information scent of the policies and reduced navigation cost
@@
All 20 participants described that the Power Meter helped them gain an intuition of the content of a policy and provided guidance. Participants described how the saturated colors, especially red and green, “piqued interest”(P4). For example, P12 described how Power Meter provided guidance and supported decision-making on which policy to pay attention to.
@@
“The color system was really guiding a lot of my decisions on which policies to read. The vibrant red was always what I was gonna try to get to 1st and the things that I cared the most about. So, finding which terms I’m agreeing to that I have the least amount of agency.” (P12)
@@
8 participants described how the document preview that appears when hovering over the Power Meter complemented and provided a more detailed overview of the policy to help them decide what to read. Additionally, 9 participants mentioned how the document preview allowed them to access information from linked pages or policies without “clicking on the link and possibly losing where I am on” (P1).
7.1.3. Summary Snippet and Color provided guidance and helped surface power and relevance within documents
18 Participants described how the color of the Summary Snippets helped surface where power lies in the contract, which is challenging to ”see” when presented in plain text with the baseline interface.
“The color coding gave me a better understanding of who’s benefiting more, how is it benefiting them, and how it affects me. Versus [in the baseline interface] it feels like it only really benefits the company itself.” (P18)
@@ 7.1.3. Summary Snippet and Color provided guidance and helped surface power and relevance within documents
While participants described focusing on the more saturated colors, especially red and green, 6 participants also described skimming the less saturated ones to validate and make sure they didn’t miss information.
7.1.4. Summary Snippet simplified and broke down dense text to support sensemaking
All participants highlighted how the Summary Snippets “broke down the entire section into a few bullet-pointed summaries” (P11), which helped them absorb the information and surface relevant pieces of information hidden in dense text. 9 participants further described how the Summary Snippets simplified the text and served as a scaffold to encourage reading the original text.
@@ 7.1.4. Summary Snippet simplified and broke down dense text to support sensemaking
Similarly, 4 participants mentioned how the highlight bars and line breaks added for each information snippet made it easier to consume the original text without having to “mentally separate the concepts” (P7). P20 described the experience of using Summary Snippets to make sense of the big ideas of the section while being able to dive deeper as forming “a web of thought and ideas: what everything kind of means together and piece by piece”. Not only do the Summary Snippets serve as an entry point to the original text, 4 participants described how the Summary Snippets helped them check their understanding after reading the original text.
7.1.5. Term definition and scenario help consume original text and envision implications
14 participants described that the underlined phrases attracted their attention, made them feel like these were phrases that “should be known and understood” (P18), and prompted them to check on their understanding of the phrase.
In addition, participants commented on how the definitions and scenarios made legal language more approachable (14/20).
@@ 7.1.5. Term definition and scenario help consume original text and envision implications
12 participants highlighted how the scenarios helped interpret the meaning and envision potential implications of the phrase in the context of the user’s intended usage of the service and was “easier to resonate”(P17).
@@
Moreover, one participant (P1) noted how the scenarios helped her come up with more questions, which she asked with the Ask function. Through this process, TermSight helped P1 to quickly obtain information from correlated policies.
7.2. RQ2: How did participants read with TermSight and its features?
@@
Figure 12. Usage of features during the reading session. Each dot represents the number of times a feature is used by one participant during a reading session. Participants used most features of TermSight. In contrast to the baseline, participants when using TermSight more frequently clicked the Summary Snippets as opposed to the section titles to navigate to the original text.
@@
7.2.1. Feature Usage
In the baseline interface, participants on average navigated to 7 policies ($\sigma=5$). Participants’ reading behaviors were similar to our findings from the formative study (§3.3.2). Participants described how they were trying to skim everything from top to bottom. When they encounter sections that were hard to read, they tended to skip. Only 11 participants clicked the section heading in the table of contents at least once to navigate to targeted sections.
As shown in Figure  12 and  13, participants used most features of TermSight throughout the session, demonstrating that TermSight provided value throughout the session. When using TermSight, participants on average navigated to 6 policies ($\sigma=3$). 19 participants used the document preview of Power Meter to gain an overview of documents in ToS for more than 1 second. 13% (17/135) of the total usage was for previewing documents hyperlinked inline, demonstrating that the Power Meter provided value also for inline hyperlinks.
Summary Snippet was the most frequently used feature, 19 participants clicked the Summary Snippet to navigate and read the original text at least once. On average, they clicked 17 Summary Snippets ($\sigma=13$). In contrast, participants rarely clicked the section headings to navigate to the original text (4 clicks in total across 3 participants). This demonstrates that Summary Snippets were potentially more useful as navigational affordances to surface and interrogate relevant information from the original text compared to section headers. We also noticed how 10 participants clicked the Highlight Bar to read the summary snippet after reading the original text, suggesting how these participants were back-checking their understanding using the summary snippets.
For Phrase Scope, 16 participants viewed the term definitions and 12 participants viewed the term scenarios. Participants on average used term definition 3 times ($\sigma=4.3$) and term scenario 1 time ($\sigma=2.7$). Out of the 77 instances when Phrase Scope was accessed, 72 were suggested by TermSight. Participants rarely manually highlighted phrases to activate phrase scope, asked questions, or checked the references provided for the definitions. While only 1 participant asked questions using Phrase Scope, P1 asked 8 questions in a 10-minute reading session.
While TermSight provided multiple reading affordances, participants were in control of deciding whether or not to rely on these features. As opposed to a consistent preference for one interaction strategy over others, we observed diverse ways that participants relied or did not rely on TermSight (Figure  14). Below, we provide case studies of these non-exclusive emergent reading strategies.
@@ 7.2.1. Feature Usage
Figure 14. Scrolling behavior of participants that employed each non-exclusive reading strategy. TermSight features two scrollable columns: Summary Snippets on the left (red lines) and Original Text on the right (blue lines). Clicking a Summary Snippet (orange X) auto-scrolls the Original Text panel to the referenced text, while clicking the Highlight Bar (light blue X) auto-scrolls the Summary panel to the corresponding Summary Snippet. (a) Summary-driven reading occurs when participants primarily read and scroll the Summary Snippets, occasionally clicking on them to interrogate the original text (19/20). (b) Original text-driven reading occurs when participants first read the original text and use the Summary Snippets as a supplement (9/20). (c) The calibration of trust mainly occurs at the beginning of the session where participants would compare and contrast the original text to the Summary Snippets with color (6/20).
7.2.2. Summary driven reading

### v2

@@ 6.4. Analysis
In this section, we introduce our analysis framework for the user study. We used a causal framework to understand the effects of the treatment and a Bayesian analysis to estimate the treatment effect (§6.4.1). Then, we discuss the thematic analysis (§6.4.2) used to analyze the qualitative responses from the semi-structured interviews.
Figure 6. The causal graph (DAG) has three types of nodes: treatment ($T$); pre-treatment covariates $D$ (demographics), $K$ (knowledge of Terms of Service); and measured outcomes $Q$ (comprehension), $R$ (recall), and $E$ (user experience). Since we use randomized assignment for the interface type (treatment vs. control), ToS type (social media vs. e-commerce), and presentation order (first vs. second), the treatment $T$ is independent of the pre-treatment covariates $D$ and $K$. To estimate the causal effect of the treatment $T$ on the outcomes $Q$, $E$, and $R$, we can use a simple regression model, without conditioning on covariates $D$ and $K$, since there is no backdoor path from the treatment $T$ to the outcome variables.
The causal graph (DAG) has three types of nodes: treatment (T) colored blue; pre-treatment covariates D (demographics), K (knowledge of Terms of Service) colored gray; and measured outcomes Q (comprehension), R (recall), and E (user experience) colored orange. Since we use randomized assignment for the interface type (treatment vs. control), ToS type (social media vs. e-commerce), and presentation order (first vs. second), the treatment T is independent of the pre-treatment covariates D and K. To estimate the causal effect of the treatment T on the outcomes Q, E, and R, we can use a simple regression model, without conditioning on covariates D and K, since there is no backdoor path from the treatment T to the outcome variables.
@@ 6.4.1. Bayesian Analysis
First, we used a structural causal framework popularized by Pearl (2009) to understand the effects of the treatment (Figure  6 shows the DAG for the study). More details about structural causal framework can be found in Appendix  F.1. Next, we used Bayesian inference to estimate the treatment effect of TermSight on the outcomes of interest. While the use of Bayesian estimates is growing in HCI (Kay et al., 2016; Koshy et al., 2023), we briefly justify its use over traditional non-Bayesian methods. First, as Kay et al. (2016) points out, a Bayesian framework leads to an accumulation of knowledge within HCI, where the posterior of the parameters in a prior work can serve as the prior in the current experiment. Second, a Bayesian model is transparent—the researcher will foreground all the assumptions in the analysis via their model. Third, use of a Bayesian framework shifts the discussion from “did it work” to “effect size of the intervention” (Kay et al., 2016). While NHST techniques can compute the confidence interval, these intervals are susceptible to misinterpretation (Belia et al., 2005; Hoekstra et al., 2014) and, importantly, underemphasized in favor of $p$-values. Finally, as  McElreath (2015) points out, a Bayesian model with the use of maximum entropy priors for the parameters (e.g., the normal distribution) is the most conservative given the evidence to estimate the effect of the treatment.
@@ 7.1. RQ1: How did participants perceive TermSight and its features?
17 participants highlighted that the baseline interface was easier to read than most of the ToS they had seen in the past, which often lacked the top navigation panel or a table of contents. When asked about which version of the interface they prefer, 19 participants preferred TermSight. One participant preferred the baseline interface, citing difficulty navigating TermSight’s two scrollable columns on a small screen with a trackpad that lacks scroll functionality. Below we present our quantitative findings on the experience outcomes (§7.1.1), followed by qualitative findings organized around recurring themes about features of TermSight (§7.1.2 – §7.1.5).
7.1.1. TermSight improved all seven user experience measures
@@
Figure  8 shows participants’ ratings of the 7 reading experience measures for both interfaces. Our Bayesian analysis (model details in Appendix F.2) reveals a significant effect of the TermSight interface on the reading experience measures. Participants found reading ToS with TermSight to require significantly less effort. When navigating ToS, deciding which policy to read and what text to read within a policy was significantly easier with TermSight. Moreover, participants were significantly more willing to read ToS and spend more time reading ToS with TermSight. The forest plots shown in Appendix Figure  14 demonstrate that the 94% Highest Posterior Density Interval (HPDI [^3] In Bayesian analysis, the 94% HPDI refers to that smallest interval of the posterior distribution that has 94% of the probability mass. It is common in Bayesian analysis to use intervals distinct (97% and 89% HPDI intervals are also used) from the typical 95% value to avoid the confusion of the frequentist confidence intervals.) for each treatment–control pair for every user experience measure does not overlap with each other. This indicates a significant positive effect of the treatment on all 7 user experience measures.
@@
Figure 7. The posterior distribution of the contrast between treatment and control on user experience is shifted to the right, with the posterior distribution of the effect size centered around 2.7 without overlapping with ROPE, suggesting a significant and large effect size.
Figure 7 shows the posterior distribution of the contrast between treatment and control on user experience. Note that the posterior distribution of the contrast is shifted to the right, with the posterior distribution of the effect size centered around 2.7 without overlapping with ROPE, suggesting a significant and large effect size.
In Figure  7, we show the contrast between the treatment and control conditions averaged across 7 user experience measures as well as the posterior of the effect size. The posterior distribution of the effect size (Cohen’s $d$) of the treatment on user experience is centered around 2.7, indicating a large effect size [^4] This is the effect size measured not on the outcome Likert scale (1–5) but on the latent scale of the model coefficients.. Because the HPDI doesn’t overlap with the ROPE [^5] We use a Region of Practical Equivalence (ROPE) of [-0.1, 0.1]. ROPE is a region where the difference between the treatment and control conditions is considered practically equivalent. This prevents us from focusing on small differences that are not practically meaningful., the effect of TermSight on user experience is significant. To further disentangle how individual features of TermSight contributed to the improvement in user experience, we present recurring themes related to participants’ perceptions and feedback on features of TermSight.
Figure 8. Participants’ ratings of reading experiences. Participants found it to be easier to find and understand relevant information with TermSight. Moreover, participants were more willing to read and spend more time on ToS with TermSight.
The figure shows seven barplots of 5-point Likert Ratings of participants’ reading experiences with TermSight and the baseline interface. Participants rated TermSight to be better for all seven experience measures. For example, participants found it to be easier to find and understand relevant information. Participants were also more willing to spend more time on the ToS with TermSight and read other ToS with TermSight.
7.1.2. Power Meter reduced navigation cost
@@
All 20 participants described that the Power Meter helped them gain an intuition of the content within each policy and supported decision-making on which policy to pay attention to.
@@
“The color system was really guiding a lot of my decisions on which policies to read. The vibrant red was always what I try to get to first and the things that I cared the most about. So, finding which terms I’m agreeing to that I have the least amount of agency.” (P12)
@@
Additionally, eight participants described how the document preview reduced navigation cost by allowing them to access information from other policies without “clicking on the link and possibly losing where I am on” (P1).
7.1.3. Summary Snippets and Color provided guidance and helped surface power and relevance within documents
Eighteen participants described how the color of the Summary Snippets helped surface where power lies within a document, which is challenging to “see” otherwise.
“The color coding gave me a better understanding of who’s benefiting more, how it’s benefiting them, and how it affects me. Versus [in the baseline], it feels like it only really benefits the company itself.” (P18)
@@ 7.1.3. Summary Snippets and Color provided guidance and helped surface power and relevance within documents
While participants described focusing on the more saturated colors, especially red and green, six participants complimented how TermSight made AI decisions (i.e., classification of power and relevance) transparent by keeping all the Summary Snippets visible and described skimming the less saturated ones to validate and make sure they didn’t miss information.
7.1.4. Summary Snippets simplified and broke down dense text to support sensemaking
Summary Snippets “broke down the entire section into a few bullet-pointed summaries” (P11), which helped participants surface and absorb relevant information hidden in dense text (20/20). Nine participants further described how the Summary Snippets simplified the text and served as a scaffold to encourage reading the original text.
@@ 7.1.4. Summary Snippets simplified and broke down dense text to support sensemaking
P20 described the experience of using Summary Snippets to make sense of the big ideas of the section while being able to dive deeper as forming “a web of thought and ideas: what everything kind of means together and piece by piece”. Not only do the Summary Snippets serve as an entry point to the original text, four participants explicitly highlighted how the Summary Snippets helped them check their understanding after reading the original text.
7.1.5. Term definitions and scenarios helped consume original text and envision implications
Fourteen participants noted that the system identified phrases attracted their attention and prompted them to check on their understanding of the phrase that they may not have noticed otherwise. Moreover, the definitions and scenarios made legal language more approachable for participants (14/20).
@@ 7.1.5. Term definitions and scenarios helped consume original text and envision implications
12 participants highlighted how the scenarios helped envision potential implications of the phrase in the context of the user’s intended usage of the service and was “easier to resonate” (P17).
@@
Moreover, one participant (P1) noted how the scenarios helped her come up with more questions, which she asked with the Ask function. Through this process, TermSight helped P1 to quickly obtain information from related policies.
7.2. RQ2: How did TermSight influence comprehension and recall?
(a) Posterior distribution of contrast (between treatment and control) on comprehension.
The figure shows the posterior distributions of the contrast between the treatment and control conditions on comprehension outcome. The posterior distribution is nearly evenly split around zero and overlaps with ROPE, indicating no effect of the treatment on the comprehension quiz questions.
(b) Posterior distribution of contrast (between treatment and control) and effect size on recall.
On the left, the figure shows the posterior distributions of the contrast between the treatment and control conditions for the recall outcomes. The posterior distribution of the treatment condition is shifted to the right, indicating a positive influence of the treatment on Recall. On the right, the figure shows the posterior distributions of the effect size of treatment centered around 0.6, suggesting a medium effect size. Yet, the HPDI has an overlap with the ROPE of [-0.1, +0.1], implying no significant effect of the treatment on the user recall.
Figure 9. In (a), the distribution centers around zero and overlaps with ROPE, indicating no effect of treatment on comprehension. (b) illustrates a positive influence of treatment on recall with the posterior distribution of the effect size centered around 0.6, suggesting a medium effect size. Because the HPDI overlaps with ROPE, there is no significant effect of treatment on recall.
In (a), the distribution centers around zero and overlaps with ROPE, indicating no effect of treatment on comprehension. (b) illustrates a positive influence of treatment on recall with the posterior distribution of the effect size centered around 0.6, suggesting a medium effect size. Because the HPDI overlaps with ROPE, there is no significant effect of treatment on recall.
7.2.1. Comprehension
Out of 6 questions, participants on average scored 2.00 ($\sigma=1.03$) when using TermSight and 2.15 ($\sigma=1.23$) when using the baseline interface. Our Bayesian analysis for the comprehension quiz (model details in Appendix  F.3) reveals no significant differences in the comprehension scores across interface conditions. As shown in Figure  9(a), the posterior distribution of the contrast between treatment and control is centered around zero. When the question ID and service type are fixed, there are significant overlaps between the 94% HPDI for each treatment–control pair as shown by the forest plots in Appendix Figure 15, suggesting no significant differences. While we aimed to design comprehension questions with one single answer based on the original text, we acknowledge that legal contracts and problems may contain room for interpretation  (Marotta-Wurgler and Stein, 2025). Therefore, we performed a supplemental, more conservative analysis by removing any question where the applicability of the original text to the question may contain room for interpretation, potentially leading to different answer choices (removed 2 out of 6 for each service type). The findings remained consistent, showing no significant difference in comprehension. More details of the analysis can be found in Appendix G.
7.2.2. Recall
Participants on average recalled 1.54 ($\sigma=1.96$) correct facts using TermSight and 1.30 ($\sigma=1.72$) correct facts using the baseline interface. Our Bayesian analysis for the recall task (model details in Appendix F.4) reveals no significant differences in the recall scores across interface conditions. Figure  9(b) shows the posterior distribution of the effect size of the treatment on recall ($\mu=0.6$, HPDI = [-0.35, 1.6]). Despite the mean being 0.6, suggesting a medium effect size, the HPDI interval has an overlap with the ROPE, implying no significant effect of the treatment on recall.
7.3. RQ3: How did participants read with TermSight and its features?
@@
Figure 10. Feature usage. Each dot represents the number of times a feature is used by one participant during a reading session. Participants used most features of TermSight. In contrast to the baseline, participants, when using TermSight, more frequently clicked the Summary Snippets as opposed to the section headers to navigate to the original text.
Box plots for the usage of features during the reading session for TermSight and the Baseline interface. Participants used most features of TermSight. In contrast to the baseline, participants, when using TermSight, more frequently clicked the Summary Snippets as opposed to the section titles to navigate to the original text.
7.3.1. Feature Usage
In the baseline interface, participants on average navigated to 7 policies ($\sigma=5$). Participants described trying to skim everything from top to bottom and often skipped difficult sections. 11 participants clicked the section heading in the table of contents at least once to navigate to targeted sections.
As shown in Figure  10 and  11, participants used most features of TermSight throughout the session. When using TermSight, participants on average navigated to 6 policies ($\sigma=3$). 19 participants hovered over a Power Meter to gain an overview of documents in ToS for more than 1 second. 13% (17/135) of the total usage was for previewing documents hyperlinked inline, demonstrating that the Power Meter provided value also for inline hyperlinks.
Summary Snippets were the most frequently used feature. 19 participants clicked the Summary Snippets to navigate and read the original text at least once ($\mu=17,\sigma=13$). In contrast, participants rarely clicked the section headers to navigate to the original text (4 clicks in total across 3 participants). This demonstrates that Summary Snippets were potentially more useful as navigational features to surface and interrogate relevant information from the original text compared to section headers. In addition, 10 participants clicked the Highlight Bar to read the Summary Snippet after reading the original text to check their understanding.
For Phrase Scope, 16 participants accessed phrase definitions and 12 participants viewed hypothetical scenarios. Out of 77 instances when Phrase Scope was accessed, 72 were suggested by TermSight. This demonstrates the value of suggesting difficult or ambiguous phrases as opposed to solely relying on users to notice them.
@@ 7.3.1. Feature Usage
Five barplots of the minute-by-minute usage of features of TermSight during the ten-minute reading task (i.e., Accessed Document Preview, Click Summary Snippet, Click Highlight Bar, Accessed Phrase Definition, and Accessed Phrase Scenario). Participants used these features throughout the reading session, rather than being limited to the beginning or end. The Summary Snippet and Document Preview are the most frequently used features throughout.
@@
Figure 12. Scrolling behavior of participants who employed each non-exclusive reading strategy. TermSight features two scrollable columns: Summary Snippets on the left (red lines) and Original Text on the right (blue lines). Clicking a Summary Snippet (orange X) auto-scrolls the Original Text panel to the referenced text, while clicking the Highlight Bar (light blue X) auto-scrolls the Summary panel to the corresponding Summary Snippet. (a) Summary-driven reading occurs when participants primarily read and scroll the Summary Snippets, occasionally clicking on them to interrogate the original text (19/20). (b) Original text-driven reading occurs when participants first read the original text and use the Summary Snippets as a supplement (9/20). (c) The calibration of trust mainly occurs at the beginning of the session, where participants would compare and contrast the original text to the Summary Snippets with color (6/20).
Visualization of the scrolling behavior of participants who employed each of the three non-exclusive reading strategies. (a) Summary-driven reading occurs when participants primarily read and scroll the Summary Snippets, occasionally clicking on them to interrogate the original text (19/20). (b) Original text-driven reading occurs when participants first read the original text and use the Summary Snippets as a supplement (9/20). (c) The calibration of trust mainly occurs at the beginning of the session where participants would compare and contrast the original text to the Summary Snippets with color (6/20).
7.3.2. Reading Strategies for Accessing Original Text
We observed diverse ways in which participants used TermSight to read ToS (Figure  12). Interestingly, these reading strategies centered around the use of the original text. In some cases, participants used TermSight’s features as entries into the original text, while in others, the features were used as verification of the original text. In still others, participants used features in tandem with the text to calibrate trust in TermSight. Below, we provide case studies of these strategies.
Summary driven reading